OpenStudy (anonymous):

What is the slope of the line that passes through (0,1) and is parallel to y = 5x + 3?

jimthompson5910 (jim_thompson5910):

what is the slope of y = 5x + 3

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

The slope of any parallel line is also 5. This is because the slopes of parallel lines are equal. So m = 5.

jimthompson5910 (jim_thompson5910):

because we know that the parallel line goes through (0,1) we know that when x = 0, y = 1

jimthompson5910 (jim_thompson5910):

so we know that m = 5 x = 0 y = 1

jimthompson5910 (jim_thompson5910):

y = mx+b y = 5x + b ... plug in the parallel slope m = 5 1 = 5*(0) + b ... plug in x = 0 and y = 1 now solve for b

OpenStudy (anonymous):

b=1 ?

jimthompson5910 (jim_thompson5910):

so m = 5 b = 1

jimthompson5910 (jim_thompson5910):

y = mx+b y = 5x + b y = 5x + 1 is your final answer